# A generalized geometric mechanics theory for multi-curve-fold origami: vertex constrained universal configurations

### v1

@@ Abstract
Folding paper along curves leads to spatial structures that have curved surfaces meeting at spatial creases, defined as curve-fold origami. In this work, we provide an Eulerian framework focusing on the mechanics of arbitrary curve-fold origami, especially for multi-curve-fold origami with vertices. We start with single-curve-fold origami that has wide panels. Wide panel leads to different domains of mechanical responses induced by various generator distributions of the curved surface. The theories are then extended to multi-curve-fold origami, involving additional geometric correlations between creases. As an illustrative example, the deformation and equilibrium configuration of origami with annular creases are studied both theoretically and numerically. Afterward, single-vertex curved origami theory is studied as a special type of multi-curve-fold origami. We find that the extra periodicity at the vertex strongly constrains the configuration space, leading to a region near the vertex that has a striking universal equilibrium configuration regardless of the mechanical properties. Both theories and numerics confirm the existence of the universality in the near-field region. In addition, the far-field deformation is obtained via energy minimization and validated by finite element analysis. Our generalized multi-curve-fold origami theory, including the vertex-contained universality, is anticipated to provide a new understanding and framework for the shape programming of the curved fold origami system.
@@ 1 Introduction
Origami is an ancient art form aiming to achieve complex spatial shapes by folding planar sheets. Classic origami only involves folding along straight lines, and it was not until the 1920s that students in Bauhaus realized sheets could be folded along curves, unveiling a novel branch that expands the traditional art form (Koschitz et al., 2008; Demaine et al., 2011). This new branch, defined as curved fold origami, opens new possibilities in origami design. Beyond art design, recently curved fold origami has found applications in soft robotics (Baek et al., 2020; Rus and Sung, 2018; Feng et al., 2024) metamaterials (Lee et al., 2021; Karami et al., 2024; Sun et al., 2024), architecture (Tachi and Epps, 2011; Mouthuy et al., 2012) and virtual reality technology (Zhang et al., 2023), due to its extraordinary abilities in energy storage, fast deployment and stiffness manipulation.
@@
The scientific exploration of curved fold origami originated with the pioneering work by David Huffman (1976). In the groundwork, curved origami structures are assumed to deform isometrically, thus they are geometrically modeled as developable surfaces meeting at space curves. Harnessing the theory of curves and surfaces in differential geometry (Do Carmo, 1976), the geometrical theories of curved origami are developed in subsequent works (Duncan and Duncan, 1982; Fuchs and Tabachnikov, 1999; Kilian et al., 2008; Demaine et al., ), revealing the geometric correlations between curved folds and bent panels. The foundational results in this area are: Given the reference strips, the two developable surfaces are determined by the space crease in the deformed configurations. The theories contribute to the computational design of developable surfaces (Rabinovich et al., 2018) and curve-fold origami (Mitani, 2011; Mitani and Igarashi, 2011; Tachi, 2013; Jiang et al., 2019; Sasaki and Mitani, 2022; Mundilova et al., 2023), as well as the predicted shape of Möbius strip (Starostin and van der Heijden, 2007; Audoly and Van der Heijden, 2023).
@@
Built upon geometric correlations, the mechanical properties of single-curved-fold origami have been investigated. The isometric deformation of curve-fold origami is determined via elastic energy (creases’ folding energy + panels’ bending energy) minimization under the constraints of developability. Mechanically, creases are modeled as elastic hinges distributed along curves and panels are modeled as unstrechable Kirchhoff plates with energy density proportional to the principle curvatures’ square. Due to the foundational results in geometric works, the bending energy of developable surfaces reduces to a $1$-dimensional integral of the energy functional associated with the geometry of reference curves. The exact form of the bending energy is derived by Wunderlich (Wunderlich, 1962; Dias and Audoly, 2015; Todres, 2015), which degenerates to the more widely used Sadowsky (1930) functional in narrow panels (Starostin and van der Heijden, 2007; Dias et al., 2012; Dias and Audoly, 2014; Yu and Hanna, 2019). Taken together, the entire elastic energy is expressed as an integral along the crease, determined by the crease’s geometry. Minimizing the energy yields the equilibrium geometry of the deformed crease as well as the configuration of the entire origami.
@@
In the area of geometric mechanics, the deformation of an annular sheet folding along its centerline is studied in Dias et al. (2012). The energy functional is expressed with curvature $\kappa$ and torsion $\tau$ of the deformed fold. Minimizing the energy leads to an equilibrium configuration that can be described analytically with $\kappa,\tau$. The theoretical results capture the buckling phenomenon well. The theory is further extended to derive a nonlinear rod model for folded elastic narrow strips (Dias and Audoly, 2014). In the work, curved origami is fitted into the framework of thin rods (Moulton et al., 2013), which simplifies numerical solving and stability analyses. In other studies, panels are constrained to deform cylindrically, thus the deformation is derived by generalized Euler-Bernoulli beam theory (Lee et al., 2018). General theory for curved origami with arbitrary deformation and reference shapes is absent. Beyond single-curve-fold origami, some works study structures with multiple curved folds (Dias and Santangelo, 2012; Liu and James, 2024). The geometric design and elastic energy of some specific multi-curve-fold origami are studied, but how geometry and elasticity contribute to the deformation of multi-curve-fold structures is still poorly understood.
@@
Based on the single-curve-fold theory, we extend the theoretical framework to multi-curve-fold origami by introducing geometric correlations between creases. Such correlations result in the propagation of generators and deformations between adjacent creases, which are solved by a numerical scheme. Accordingly, the energy distribution is also derived, leading to the geometric mechanics theory of multi-curve-fold origami. Validations are made on an illustrative structure with multiple annular folds, where theoretical and finite element analysis (FEA) results fit well.
More interestingly, we find that a mechanics-independent deformed configuration emerges when the multiple creases meet at a vertex. We prove that in the system of multi-curve-fold origami with a vertex, the periodicity at the vertex yields a universal equilibrium configuration in the domain near the vertex, regardless of mechanical properties (stiffness of folds and panels) and geodesic curvatures of folds. Specifically, the deformed configuration is determined only by the folding angle at the vertex. In contrast, in the far-field domain determined by the aforementioned geometric correlations, the energy minimization yields various equilibrium configurations for different mechanical constants. Both results are validated by FEA. The theories allow us to predict the deformation of multi-curve-fold origami with vertices.
@@ 2.1 Preliminaries: geometry of single-curve-fold origami
We start with the classic differential geometry for the shape description of curved fold origami in the deformed domain, i.e., in the Eulerian framework. In this paper, we assume the origami structure is deformed isometrically from a flat sheet. Therefore, the flanks are modeled as developable surfaces, which are generically given by
@@
Here $\mathbf{r}_{0}(S)$ is a selected reference curve on the surface, and $\mathbf{l}(S)$ is the generator at the arclength $S$. In curved-fold origami, it is convenient to choose the curved crease as the reference curve and the flanks on the two sides are described by Eq. (1). Solving these two equations yields strong geometric correlations between the reference curve and the developable surface, which are given explicitly via solving the kinematic equations of Frenet frames and Darboux frames (shown in Fig. 1(a)).
@@
where $\kappa$ and $\tau$ are the curvature and torsion, which determine the shape of the space curve uniquely up to rigid motions by the fundamental theorem of curves (Do Carmo (1976)).
@@
where $\omega_{ij}$ is the differential form representing the rotation of the Darboux frame. The nonzero principle curvature $\kappa_{m}$ satisfies
@@
$\kappa_{m}=\mathrm{Tr}(B).$
Geodesic curvature $\kappa_{g}$ and normal curvature $\kappa_{n}$ of the curve on surface are defined as
@@
and the generator vector $\mathbf{l}$ is selected as
@@
and the nonzero principle curvature is derived by
@@
On the premise of developability, the generators cannot intersect with each other inside the surface, thus $(\partial\mathbf{r}(S,v)/\partial S)\cdot\mathbf{e}_{1}=1+vt^{\prime}-v\kappa_{g}-vt^{2}\kappa_{g}$ stays positive along any generator. The generators may intersect at boundaries of the surface, forming a conical point introducing singularity of $\kappa_{m}$.
@@
Therefore, the principle curvature $\kappa_{m}(S,v)$ distribution on the panel depends on the curvature $\kappa(S)$ and torsion $\tau(S)$ of the curved fold, according to Eq. (7), (11) and (12). Symmetrically we derive the geometric variables in the opposite panel
@@
$\kappa_{m\pm}$ $=\frac{(1+t_{\pm}^{2})\kappa_{n{\pm}}}{1\pm v_{\pm}t_{\pm}^{\prime}-v_{\pm}\kappa_{g{\pm}}-v_{\pm}t_{\pm}^{2}\kappa_{g\pm}},$
@@
$\phi=\left\langle\mathbf{e}_{3+},\mathbf{e}_{3-}\right\rangle=|\arccos(\frac{\kappa_{n+}}{\kappa})\pm\arccos(\frac{\kappa_{n-}}{\kappa})|,$ (14)
@@
2.2 Domain classification for curved fold origami with wide flanks
@@
Crease-crease domain. In this domain, the generators start from one crease and end at the same crease. At a limiting point $S_{c}^{*}$, the length of the generator tends to zero, where $\kappa_{g+}>0$ and $t_{+}$ jumps rapidly from +$\infty$+ ∞ to $\infty$- ∞. In the following discussion we discuss how generator distribute in panel --. To derive the generator distribution function $t_{-}(S)$, we focus on an infinitesimal region around $(S^{*}_{c})$ and define $\kappa^{*}_{n}(\xi)=\kappa_{n+}(S^{*}_{c}+\xi)$, $\kappa^{*}_{g}(\xi)=\kappa_{g+}(S^{*}_{c}+\xi)$, $\kappa^{*}(\xi)=\kappa(S^{*}_{c}+\xi)$, $\tau^{*}(\xi)=\tau(S^{*}_{c}+\xi)$ and $t^{*}_{\pm}(\xi)=t_{\pm}(S^{*}_{c}+\xi)$ with $\xi\rightarrow 0$. Due to possible singularities in Eq. (15), $t_{-}(S)$ is given by analyzing the the order of the parts $\kappa_{g}^{*}\kappa_{n}^{*\prime}/\kappa^{*2}\kappa_{n}^{*}$, $\kappa_{g}^{*\prime}/\kappa^{*2}$ and $\tau^{*}/\kappa_{n}^{*}$ in Eq. (15).
@@
To keep the principle curvature $\kappa_{m}$ finite, an infinite $t$ derived from Eq. (17) should keep $(1+t^{*2}(\xi))\kappa^{*}_{n}(\xi)$ finite. Therefore, $\kappa^{*}_{n}(\xi)$ is at most $O(\xi^{2}\kappa^{*}_{g}(\xi))$. We then derive
@@
Therefore, $t$ in panel -- jumps from $\infty$- ∞ to +$\infty$+ ∞ rapidly. To avoid the intersection of generators in panel --, panel -- must be divided into two pieces at $S_{c}^{*}$. To help understand this conclusion we give a simple example of curved origami with crease-crease domain. As demonstrated in Fig. 2(b), a semi-circle fold is folded into
@@ 2.3.1 Energy distribution in single-curve-fold origami
The elastic energy of curved fold origami consists of folding energy concentrated on creases and bending energy distributed in panels. For folding energy, we assume that folds are modeled as linear elastic torsional springs with direction $\mathbf{t}(S)$. The energy of the fold between $(S,S+dS)$ is expressed as
@@
where $S_{c-c},S_{c-e}$ and $S_{e-e}$ represent the arclength parameter of the crease-crease, crease-edge, and edge-edge domain, respectively. Taking together the integral of $\varepsilon_{b}(S)$ along the crease and the edge-edge segment yields the bending energy of the whole panel. Therefore, the total elastic energy of the curved fold origami is

### v2

@@ Abstract
Folding paper along curves leads to spatial structures that have curved surfaces meeting at spatial creases, defined as curve-fold origami. In this work, we provide an Eulerian framework focusing on the mechanics of arbitrary curve-fold origami, especially for multi-curve-fold origami with vertices. We start with single-curve-fold origami that has wide panels. Wide panel leads to different domains of mechanical responses induced by various generator distributions of the curved surface. The theories are then extended to multi-curve-fold origami, involving additional geometric correlations between creases. As an illustrative example, the deformation and equilibrium configuration of origami with annular creases are studied both theoretically and numerically. Afterward, single-vertex curved origami theory is studied as a special type of multi-curve-fold origami. We find that the extra periodicity at the vertex strongly constrains the configuration space, leading to a region near the vertex that has a striking universal equilibrium configuration regardless of the mechanical properties. Both theories and numerics confirm the existence of the universality in the near-field region. In addition, the far-field deformation is obtained via energy minimization and validated by finite element analysis. Our generalized multi-curve-fold origami theory, including the vertex-contained universality, is anticipated to provide a new understanding and framework for the shape programming of the curve-fold origami system.
@@ 1 Introduction
Origami is an ancient art form aiming to achieve complex spatial shapes by folding planar sheets. Classic origami only involves folding along straight lines, and it was not until the 1920s that students in Bauhaus realized sheets could be folded along curves, unveiling a novel branch that expands the traditional art form (Koschitz et al., 2008; Demaine et al., 2011). This new branch, defined as curve-fold origami, opens new possibilities in origami design. Beyond art design, recently curve-fold origami has found applications in soft robotics (Baek et al., 2020; Rus and Sung, 2018; Feng et al., 2024), metamaterials (Lee et al., 2021; Karami et al., 2024; Sun et al., 2024), architecture (Tachi and Epps, 2011; Mouthuy et al., 2012) and virtual reality technology (Zhang et al., 2023), due to its extraordinary abilities in energy storage, fast deployment and stiffness manipulation.
@@
The scientific exploration of curve-fold origami originated with the pioneering work by David Huffman (1976). In the groundwork, curved origami structures are assumed to deform isometrically, thus they are geometrically modeled as developable surfaces meeting at space curves. Harnessing the theory of curves and surfaces in differential geometry (Do Carmo, 1976), the geometrical theories of curved origami are developed in subsequent works (Duncan and Duncan, 1982; Fuchs and Tabachnikov, 1999; Kilian et al., 2008; Demaine et al., ), revealing the geometric correlations between curved folds and bent panels. The foundational results in this area are: Given the reference strips, the two developable surfaces are determined by the space crease in the deformed configurations. The theories contribute to the computational design of developable surfaces (Rabinovich et al., 2018) and curve-fold origami (Mitani, 2011; Mitani and Igarashi, 2011; Tachi, 2013; Jiang et al., 2019; Sasaki and Mitani, 2022; Mundilova et al., 2023), as well as the predicted shape of Möbius strip (Starostin and van der Heijden, 2007; Audoly and Van der Heijden, 2023).
@@
Built upon geometric correlations, the mechanical properties of single-curve-fold origami have been investigated. The isometric deformation of curve-fold origami is determined via elastic energy (creases’ folding energy + panels’ bending energy) minimization under the constraints of developability. Mechanically, creases are modeled as elastic hinges distributed along curves and panels are modeled as unstrechable Kirchhoff plates with energy density proportional to the principal curvatures’ square. Due to the foundational results in geometric works, the bending energy of developable surfaces reduces to a $1$-dimensional integral of the energy functional associated with the geometry of reference curves. The exact form of the bending energy is derived by Wunderlich (Wunderlich, 1962; Dias and Audoly, 2015; Todres, 2015), which degenerates to the more widely used Sadowsky (1930) functional in narrow panels (Starostin and van der Heijden, 2007; Dias et al., 2012; Dias and Audoly, 2014; Yu and Hanna, 2019). Taken together, the entire elastic energy is expressed as an integral along the crease, determined by the crease’s geometry. Minimizing the energy yields the equilibrium geometry of the deformed crease as well as the configuration of the entire origami.
@@
In the area of geometric mechanics, the deformation of an annular sheet folding along its centerline is studied in Dias et al. (2012). The energy functional is expressed with curvature $\kappa$ and torsion $\tau$ of the deformed fold. Minimizing the energy leads to an equilibrium configuration that can be described analytically with $\kappa,\tau$. The theoretical results capture the buckling phenomenon well. The theory is further extended to derive a nonlinear rod model for folded elastic narrow strips (Dias and Audoly, 2014). In the work, curved origami is fitted into the framework of thin rods (Moulton et al., 2013), which simplifies numerical solving and stability analyses. In other studies, panels are constrained to deform cylindrically, thus the deformation is derived by generalized Euler-Bernoulli beam theory (Lee et al., 2018). However, a general theory for curved origami with arbitrary deformation and reference shapes is absent. Beyond single-curve-fold origami, some previous works have studied multiple curved folds, including the geometric design (Dias, 2012; Liu and James, 2024), and the energy of folds with constant curvature and torsion (for example, a comprehensive study for helicoids) (Dias and Santangelo, 2012). However, the Euler-Lagrange equations (or equilibrium equations) for multi-curve-fold systems remain unknown. Thus, how geometry and elasticity contribute to the deformation is still poorly understood. Furthermore, the theory for multi-curve-fold origami with vertices is still lacking.
@@
Based on the single-curve-fold theory, we extend the theoretical framework to multi-curve-fold origami by introducing geometric correlations between creases. Such correlations result in the propagation of generators and deformations between adjacent creases, which are solved by a numerical scheme. Accordingly, the energy distribution is derived, leading to the geometric mechanics theory of multi-curve-fold origami. Beyond existing works, we derive the Euler-Lagrange equations revealing how geometry and elasticity determine the multi-curve-fold systems’ deformation. Validations are made on an illustrative structure with multiple annular folds, where theoretical and finite element analysis (FEA) results fit well.
More interestingly, we find that a mechanics-independent deformed configuration emerges when the multiple creases meet at a vertex. We prove that in the system of multi-curve-fold origami with a vertex, the periodicity at the vertex yields a universal equilibrium configuration in the domain near the vertex, regardless of mechanical properties (stiffness of folds and panels) and geodesic curvatures of folds. Specifically, the deformed configuration is determined only by the folding angle at the vertex. In contrast, in the far-field domain determined by the aforementioned geometric correlations, the energy minimization yields various equilibrium configurations for different mechanical constants. Both results are validated by FEA. The theories pioneerly indicate how multi-curve-fold origami with vertices deform.
@@ 2.1 Preliminaries: geometry of single-curve-fold origami
We start with the classic differential geometry for the shape description of curve-fold origami in the deformed domain, i.e., in the Eulerian framework. In this paper, we assume the origami structure is deformed isometrically from a flat sheet. Therefore, the flanks are modeled as developable surfaces, which are generically given by
@@
Here $\mathbf{r}_{0}(S)$ is a selected reference curve on the surface, and $\mathbf{l}(S)$ is the generator at the arclength $S$. In curve-fold origami, it is convenient to choose the curved crease as the reference curve and the flanks on the two sides are described by Eq. (1). Solving these two equations yields strong geometric correlations between the reference curve and the developable surface, which are given explicitly via solving the kinematic equations of Frenet frames and Darboux frames (shown in Fig. 1(a)).
@@
where $\kappa$ and $\tau$ are the curvature and torsion, which determine the shape of the space curve uniquely up to rigid motions by the fundamental theorem of curves (Do Carmo, 1976).
@@
where $\omega_{ij}$ is the differential form representing the rotation of the Darboux frame. According to classical differential geometry (page 63 of (Do Carmo, 1976)), the nonzero principal curvature $\kappa_{m}$ satisfies
@@
$\kappa_{m}=\mathrm{Tr}(B),$
where $[\omega_{13},\omega_{23}]$ and $[\omega_{1},\omega_{2}]$ are $1\times 2$ matrices and $B$ is a $2\times 2$ symmetric matrix related to the second fundamental form of the surface. Geodesic curvature $\kappa_{g}$ and normal curvature $\kappa_{n}$ of the curve on surface are defined as
@@
where $\left\langle\mathbf{e}_{1},\mathbf{l}\right\rangle$ represents the angle between the vectors $\mathbf{e}_{1}$ and $\mathbf{l}$, ranging from 00 to $\pi$. The generator vector $\mathbf{l}$ is selected as
@@
and the nonzero principal curvature is derived by
@@
On the premise of developability, the generators cannot intersect with each other inside the surface, otherwise, the intersection will introduce divergent principal curvature and divergent bending energy discussed later. To see this, let $\mathbf{r}(S^{\star},v^{\star})$ be the generator intersection, satisfying $(\partial\mathbf{r}(S^{\star},v^{\star})/\partial S)\cdot\mathbf{e}_{1}=0$. By direct calculation, we have $(\partial\mathbf{r}(S^{\star},v^{\star})/\partial S)\cdot\mathbf{e}_{1}=1+v^{\star}t^{\prime}-v^{\star}\kappa_{g}-v^{\star}t^{2}\kappa_{g}=0$, resulting in divergent principal curvature at $(S^{\star},v^{\star})$ according to Eq. (11).
@@
Therefore, the principal curvature $\kappa_{m}(S,v)$ distribution on the panel depends on the curvature $\kappa(S)$ and torsion $\tau(S)$ of the curved fold, according to Eq. (7), (11) and (12). Symmetrically we derive the geometric variables in the opposite panel
@@
$\kappa_{m_{\pm}}$ $=\frac{(1+t_{\pm}^{2})\kappa_{n\pm}}{1\pm v_{\pm}t_{\pm}^{\prime}-v_{\pm}\kappa_{g\pm}-v_{\pm}t_{\pm}^{2}\kappa_{g\pm}},$
@@
$\phi=\langle\mathbf{e}_{3+},\mathbf{e}_{3-}\rangle=|\arccos(\frac{\kappa_{n+}}{\kappa})\pm\arccos(\frac{\kappa_{n-}}{\kappa})|,$ (14)
@@
2.2 Domain classification for curve-fold origami with wide flanks
@@
Crease-crease domain. In this domain, the generators start from one crease and end at the same crease. At a limiting point $S_{c}^{*}$, the length of the generator tends to zero, where $\kappa_{g+}>0$ and $t_{+}$ jumps rapidly from +$\infty$+ ∞ to $\infty$- ∞. In the following discussion we discuss how generator distribute in panel --. To derive the generator distribution function $t_{-}(S)$, we focus on an infinitesimal region around $(S^{*}_{c})$ and define $\kappa^{*}_{n}(\xi)=\kappa_{n+}(S^{*}_{c}+\xi)$, $\kappa^{*}_{g}(\xi)=\kappa_{g+}(S^{*}_{c}+\xi)$, $\kappa^{*}(\xi)=\kappa(S^{*}_{c}+\xi)$, $\tau^{*}(\xi)=\tau(S^{*}_{c}+\xi)$ and $t^{*}_{\pm}(\xi)=t_{\pm}(S^{*}_{c}+\xi)$ with $\xi\rightarrow 0$. Due to possible singularities in Eq. (15), $t_{-}(S)$ is given by analyzing the order of the parts $\kappa_{g}^{*}\kappa_{n}^{*\prime}/\kappa^{*2}\kappa_{n}^{*}$, $\kappa_{g}^{*\prime}/\kappa^{*2}$ and $\tau^{*}/\kappa_{n}^{*}$ in Eq. (15).
@@
To keep the principal curvature $\kappa_{m}$ finite, an infinite $t$ derived from Eq. (17) should keep $(1+t^{*2}(\xi))\kappa^{*}_{n}(\xi)$ finite. Therefore, $\kappa^{*}_{n}(\xi)$ is at most $O(\xi^{2}\kappa^{*}_{g}(\xi))$. We then derive
@@
Therefore, $t-$ in panel -- jumps from $\infty$- ∞ to +$\infty$+ ∞ rapidly. To avoid the intersection of generators in panel --, panel -- must be divided into two pieces at $S_{c}^{*}$. To help understand this conclusion we give a simple example of curved origami with crease-crease domain. As demonstrated in Fig. 2(b), a semi-circle fold is folded into
@@ 2.3.1 Energy distribution in single-curve-fold origami
The elastic energy of curve-fold origami consists of folding energy concentrated on creases and bending energy distributed in panels. For folding energy, we assume that folds are modeled as linear elastic torsional springs with direction $\mathbf{t}(S)$. The energy of the fold between $(S,S+dS)$ is expressed as
@@
where $S_{c-c},S_{c-e}$ and $S_{e-e}$ represent the arclength parameter of the crease-crease, crease-edge, and edge-edge domain, respectively. Taking together the integral of $\varepsilon_{b}(S)$ along the crease and the edge-edge segment yields the bending energy of the whole panel. Therefore, the total elastic energy of the curve-fold origami is